OpenStudy (anonymous):

how do you simplify (x^2 - 9)(2x) - (x^2)(2x) to get -18x? This is in several problems in my book but the steps are never explained.

OpenStudy (klimenkov):

\[(x^2-9)\cdot2x-x^2\cdot2x=(x^2-9-x^2)\cdot2x=(x^2-x^2-9)\cdot2x=-9\cdot2x=-18x\]

OpenStudy (klimenkov):

The last is \[-18x\]

OpenStudy (anonymous):

Thanks!

OpenStudy (klimenkov):

Ask if you can not understand something.